# Semiclassical solution of black hole information paradox

### v1

@@ Abstract
We resolve black hole information paradox within semiclassical gravity, in a manner that does not depend on details of unknown quantum gravity. Our crucial insight is that outgoing Hawking particles are physical only at distances much larger than the Schwarzschild radius, so they are created far from the horizon and entangled with degrees of freedom at smaller distances. The later degrees of freedom can be understood as quasi-classical coherent states, implying that Hawking radiation is accompanied by additional radiation similar to classical radiation by which the black hole loses hair during the classical gravitational collapse. The two kinds of radiation are entangled, which resolves black hole information paradox.
@@
One of the central questions that we deal with is where are the outgoing Hawking particles created unruh_where1 ; unruh_where2 ; giddings_where ; hod_where ? The usual picture is that they are created near the horizon. By contrast, we argue that they are created far from the horizon, at a distance much larger than the Schwarzschild radius $2M$ (with $M$ being the mass of the black hole, and we work in units $\hbar=c=G_{N}=k_{B}=1$). Our argument for this claim is based on considerations of foundations of quantum mechanics itself laloe . The contextuality theorems of quantum mechanics KS ; bell ; peres ; laloe show that it is not consistent to assume that measurements just reveal the values of observables that the quantum system possessed before measurements. Instead, it is the measurement that creates physical reality as we perceive it. Of course, the measurement does not need to be performed by a human made apparatus. Instead, the measurement of a quantum subsystem is associated with decoherence decoh1 ; decoh2 caused by its environment, which makes the density matrix of a subsystem diagonal in a certain “preferred” basis determined by interaction with the environment. Applying this logic to Hawking radiation, it follows that it is inconsistent to speak of outgoing Hawking particles before they are measured by the environment. The outgoing Hawking particles are defined as positive frequency modes with respect to the Schwarzschild time coordinate bd , which is a physical time for observers static with respect to the black hole. Such static observers are approximately inertial when they are far from the horizon, namely at $r\gg 2M$. But at smaller distances such observers are non-inertial, so the corresponding “particles” can only by physical if the environment that measures the particles is itself static and thus non-inertial. In principle it is possible to have such a measuring apparatus, but then it measures the particles by the Unruh effect unruh , which requires its own source of energy (such as the fuel of the rocket that keeps a detector hovering at a static position with respect to the black hole), so such a physical realization of Hawking particles at smaller distances is compensated by the loss of energy of this source mochizuki , which does not need to decrease the black hole mass. Thus only particles at $r\gg 2M$ are created by measurement without any additional source of energy for the detector, which makes particles at $r\gg 2M$ much more typical than those at smaller $r$. Hence, for practical purposes, the creation of Hawking particles is naturally interpreted as something that typically happens far from the horizon, at $r\gg 2M$.
@@
The rest of the paper is organized as follows. To build intuition for our approach to black hole information paradox, we first consider two simple analogies, the hydrogen atom information “paradox”, and the case of two entangled harmonic oscillators, in settings that resemble our view of Hawking radiation. Then we consider the quantum state of matter in a classical black hole background and rewrite it in terms of physical Hawking particle states at large distances from the horizon, entangled with degrees of freedom at smaller distances. We rewrite the state of the latter degrees of freedom in the basis of quasi-classical coherent states, and argue that radiation of Hawking particles is accompanied with an additional quasi-classical radiation, the latter being similar to radiation produced during the classical gravitational collapse by which a classical black hole loses classical hair. We also discuss the role of more exotic quantum gravitational objects, proposed in the literature as possible solutions of black hole information paradox.
@@
Introducing the unit operator in the position basis $1=\int d^{3}x\,|{\bf x}\rangle\langle{\bf x}|$, the initial ground state can be written as
@@
describing a perfect correlation between positions of the two particles. The particle-$B$ is in the mixed state
@@
Next suppose that the interaction between the two particles is turned off. The two particles will remain entangled. However, the electron (particle-$A$) will be then captured by the proton again, so it will end up in the ground state $|\psi_{0}\rangle$. The ground state is unique, so its entropy is zero. The particle-$B$, on the other hand, remains in the mixed state with non-zero entropy and it cannot longer be correlated with the particle-$A$, because the latter has zero entropy. This establishes the hydrogen atom information “paradox”.
The solution of the “paradox”, of course, is that the electron radiates again, by the same mechanism as the first time. This secondary radiation is entangled with particle-$B$, which resolves the “paradox”. We shall argue that the black hole information paradox resolves in an analogous way, namely that, in addition to the outgoing Hawking particles (analogous to particle-$B$) there is also a secondary radiation of gravitational waves (analogous to secondary electromagnetic radiation), so that the outgoing Hawking particles are correlated with the secondary radiation.
@@
$|n\rangle=\frac{a^{\dagger}|0\rangle}{\sqrt{n!}}.$ (4)
Suppose that they are in the entangled state of the thermal form
@@
where $c_{n}(z)={\cal N}e^{-\beta\omega n/2}{}_{A}\langle z|n\rangle_{A}$. The $|c_{n}(z)|^{2}$ is the probability density that the oscillator-$B$ will be found in the quantum state $|n\rangle_{B}$, and the oscillator-$A$ in the quasi-classical coherent state $|z\rangle_{A}$.
Semiclassical black hole.—Consider a black hole of mass $M$, described either by Schwarzschild $S$-coordinates with coordinate singularity at the horizon at $r=2M$, or by Kruskal $K$-coordinates without a coordinate singularity. The initial state of the quantum field $\phi(x)$ in the classical black hole background is the Kruskal vacuum $|O_{K}\rangle$ bd (known also as Hartle-Hawking vacuum), naturally associated with the $K$-coordinates. It is related to the $S$-vacuum $|O_{S}\rangle$, naturally associated with $S$-coordinates, through the formula bd
@@
At $r\gg 2M$, the “particles” created by $b_{2k}^{\dagger}$ behave as ordinary particles in quantum field theory in flat spacetime. For that reason, those “particles” are considered in the literature to be the actual physical particles, so (9) is interpreted as physical particle creation in zone-2. However, the black hole interior in zone-1 has a limited number of degrees of freedom, i.e. limited entropy, and it turns out that it does not have enough entropy to explain the entropy of particles created in zone-2, through the entanglement between zone-2 and zone-1. This is the essence of black hole information paradox gid ; math1 ; math2 ; hoss ; dundar ; harlow ; polchinski ; chakra ; marolf ; fabbri .
Our crucial new insight is the following. As we already explained in Introduction, the particles created by $b_{2k}^{\dagger}$ are physical at very large distances $r\gg 2M$, where the spacetime is essentially Minkowski spacetime, but they are not physical for intermediate values of $r$. Hence we split zone-2 into two zones, called zone-$A$ and zone-$B$, defined as follows
@@
where $R^{\prime}$ is some fixed large radius $R^{\prime}\gg 2M$. Hence, as shown by (35) in Appendix, the creation operator $b_{2k}^{\dagger}$ can be decomposed into two operators
@@
The upshot of this calculation is Eq. (20), which shows that physical particles in zone-$B$ are entangled with the states in zone-$(1\!\cup\!A)$, the union of zone-1 and zone-$A$. Since $R^{\prime}\gg 2M$, zone-$(1\!\cup\!A)$ is much larger than the black hole interior zone-1. Moreover, most of the zone-$(1\!\cup\!A)$ is outside of the black hole. Thus the zone-$(1\!\cup\!A)$ contains a lot of entropy, and most of it is not hidden behind the horizon. Since the “particles” in zone-$(1\!\cup\!A)$ are not physical particles, the physical particle creation can be considered to occur at the sphere of radius $R^{\prime}\gg 2M$. The entanglement entropy of particles created at $R^{\prime}$ can easily be explained through entanglement with degrees of freedom in zone-$(1\!\cup\!A)$, which, in principle, resolves the black hole information paradox.
@@
Then we assume that, in zone-$(1\!\cup\!A)$, there is an over-complete basis of quasi-classical coherent states $|Z_{1\cup A}\rangle$ for both gravitational and matter fields, $|Z_{1\cup A}\rangle=|Z\rangle_{\rm gravitational}|Z\rangle_{\rm matter}$, obeying
@@
Without a full theory of quantum gravity, we cannot specify the coherent states $|Z_{1\cup A}\rangle$ explicitly. Nevertheless, their essential properties can easily be inferred by heuristic arguments. It is natural to expect that a typical coherent state resembles a classical configuration of gravitational and matter fields. Hence, a typical coherent state behaves approximately classically. In other words, it behaves as a classical hair in zone-$(1\!\cup\!A)$. But we know from classical no hair theorem hawking-ellis ; frolov-novikov that classical hair soon gets radiated away, in terms of gravitational (and matter) waves. Thus our approach predicts that there are two kinds of radiation from the black hole. First there are Hawking particles created in zone-$B$, and second there is a quasi-classical radiation of gravitational and matter waves created in zone-$A$. The two kinds of radiation are entangled with each other, which resolves the black hole information paradox.
@@
Such a resolution of the paradox does not need depend on any details of quantum gravity and does not involve any exotic phenomena that would contradict general expectations from semiclassical gravity. In particular, a freely falling observer near the horizon will not observe anything out of the ordinary. Indeed, since Hawking particles are not physical near the horizon, such a freely falling observer near the horizon will not observe any Hawking particles.
@@
Appendix A Bogoliubov coefficients
@@
are the Bogoliubov coefficients. For the same reason, the scalar products $(f_{2k},f_{Ak}^{*})$ and $(f_{2k},f_{Bk}^{*})$, i.e., the $\beta$-coefficients in (A), are also negligible, so (33) reduces to

### v2

@@ Abstract
We resolve black hole information paradox within semiclassical gravity, in a manner that does not depend on details of unknown quantum gravity. Our crucial insight is that outgoing Hawking particles are physical only far from the black hole horizon, so they are created far from the horizon and entangled with degrees of freedom closer to the horizon. The later degrees of freedom can be understood as quasi-classical coherent states, implying that Hawking radiation is accompanied with additional radiation similar to classical radiation by which the black hole loses hair during the classical gravitational collapse. The two kinds of radiation are entangled, which resolves black hole information paradox.
@@
One of the central questions that we deal with is where are the outgoing Hawking particles created unruh_where1 ; unruh_where2 ; giddings_where ; hod_where ? The usual picture is that they are created near the horizon. By contrast, we argue that they are created far from the horizon, at a distance much larger than the Schwarzschild radius $2M$ (with $M$ being the mass of the black hole, and we work in units $\hbar=c=G_{N}=k_{B}=1$). Our argument for this claim is based on considerations of foundations of quantum mechanics itself laloe . The contextuality theorems of quantum mechanics KS ; bell ; peres ; laloe show that it is not consistent to assume that measurements just reveal the values of observables that the quantum system possessed before measurements. Instead, it is the measurement that creates physical reality as we perceive it. Of course, the measurement does not need to be performed by a human made apparatus. Instead, the measurement of a quantum subsystem is associated with decoherence decoh1 ; decoh2 caused by its environment, which makes the density matrix of a subsystem diagonal in a certain “preferred” basis determined by interaction with the environment. Applying this logic to Hawking radiation, it follows that it is inconsistent to speak of outgoing Hawking particles before they are measured by the environment. The outgoing Hawking particles are defined as positive frequency modes with respect to the Schwarzschild time coordinate bd , which is a physical time for observers static with respect to the black hole. Such static observers are approximately inertial when they are far from the horizon, namely at $r\gg 2M$. But at smaller distances such observers are non-inertial, so the corresponding “particles” can only be physical if the environment that measures the particles is itself static and thus non-inertial. In principle it is possible to have such a measuring apparatus, but then it measures the particles by the Unruh effect unruh , which requires its own source of energy (such as the fuel of the rocket that keeps a detector hovering at a static position with respect to the black hole), so such a physical realization of Hawking particles at smaller distances is compensated by the loss of energy of this source mochizuki , which does not need to decrease the black hole mass. Thus only particles at $r\gg 2M$ are created by measurement without any additional source of energy for the detector, which makes particles at $r\gg 2M$ much more typical than those at smaller $r$. Hence, for practical purposes, the creation of Hawking particles is naturally interpreted as something that typically happens far from the horizon, at $r\gg 2M$.
@@
The rest of the paper is organized as follows. To build intuition for our approach to black hole information paradox, we first consider two simple analogies, the hydrogen atom information “paradox”, and the case of two entangled harmonic oscillators, in settings that resemble our view of Hawking radiation. Then we consider the quantum state of matter in a classical black hole background and rewrite it in terms of physical Hawking particle states at large distances from the horizon, entangled with degrees of freedom at smaller distances. We rewrite the state of the latter degrees of freedom in the basis of quasi-classical coherent states, and argue that radiation carried by Hawking particles is accompanied with an additional quasi-classical radiation, the latter being similar to radiation produced during the classical gravitational collapse by which a classical black hole loses classical hair. We also discuss the role of more exotic quantum gravitational objects, proposed in the literature as possible solutions of black hole information paradox.
@@
Introducing the unit operator in the position basis $1=\int d^{3}x\,|{\bf x}\rangle\langle{\bf x}|$, the ground state can be written as
@@
describing a perfect correlation between positions of the two particles. Particle-$B$ is in the mixed state
@@
Next suppose that the interaction between the two particles is turned off. The two particles will remain entangled. However, the electron (particle-$A$) will be then captured by the proton again, so it will end up in the ground state $|\psi_{0}\rangle$. The ground state is unique, so its entropy is zero. Particle-$B$, on the other hand, remains in the mixed state with non-zero entropy and it cannot longer be correlated with particle-$A$, because the latter has zero entropy. This establishes the hydrogen atom information “paradox”.
The solution of the “paradox”, of course, is that the electron radiates again (see Appendix A), by the same mechanism as the first time. This secondary radiation is entangled with particle-$B$, which resolves the “paradox”. We shall argue that the black hole information paradox resolves in an analogous way, namely that, in addition to the outgoing Hawking particles (analogous to particle-$B$) there is also a secondary radiation of gravitational waves (analogous to secondary electromagnetic radiation), so that the outgoing Hawking particles are correlated with the secondary radiation.
@@
$|n\rangle=\frac{(a^{\dagger})^{n}|0\rangle}{\sqrt{n!}}.$ (4)
Suppose that the two oscillators are in the entangled state of the thermal form
@@
where $c_{n}(z)={\cal N}e^{-\beta\omega n/2}{}_{A}\langle z|n\rangle_{A}$. The $|c_{n}(z)|^{2}$ is the probability density that oscillator-$B$ will be found in the quantum state $|n\rangle_{B}$, and oscillator-$A$ in the quasi-classical coherent state $|z\rangle_{A}$.
Semiclassical black hole.—Consider a black hole of mass $M$, described either by Schwarzschild $S$-coordinates with coordinate singularity at the horizon at $r=2M$ (where $r$ is the usual Schwarzschild radial coordinate), or by Kruskal $K$-coordinates without a coordinate singularity. The initial state of the quantum field $\phi(x)$ in the classical black hole background is the Kruskal vacuum $|O_{K}\rangle$ bd (known also as Hartle-Hawking vacuum), naturally associated with $K$-coordinates. It is related to the $S$-vacuum $|O_{S}\rangle$, naturally associated with $S$-coordinates, through the formula bd
@@
At $r\gg 2M$, the “particles” created by $b_{2k}^{\dagger}$ behave as ordinary particles in quantum field theory in flat Minkowski spacetime. For that reason, those “particles” are considered in the literature to be the actual physical particles, so (9) is interpreted as physical particle creation in zone-2. However, the black hole interior in zone-1 has a limited number of degrees of freedom, i.e. limited entropy, and it turns out that it does not have enough entropy to explain the entropy of particles created in zone-2, through the entanglement between zone-2 and zone-1. This is the essence of black hole information paradox gid ; math1 ; math2 ; hoss ; dundar ; harlow ; polchinski ; chakra ; marolf ; fabbri .
Our crucial new insight is the following. As we already explained in Introduction, the particles created by $b_{2k}^{\dagger}$ are physical at very large distances $r\gg 2M$, where the spacetime is essentially Minkowski spacetime, but they are not physical at intermediate distances $r$. Hence we split zone-2 into two zones, called zone-$A$ and zone-$B$, defined as follows:
@@
where $R^{\prime}$ is some fixed large radius $R^{\prime}\gg 2M$. (The precise value of $R^{\prime}$ is not important. The two zones with a sharp boundary serve as a simplified model, while in a more realistic model there would be no sharp boundary between the two zones.) Hence, as shown by (37) in Appendix B, the creation operator $b_{2k}^{\dagger}$ can be decomposed into two operators
@@
The upshot of this calculation is Eq. (20), which shows that physical particles in zone-$B$ are entangled with the states in zone-$(1\!\cup\!A)$, the union of zone-1 and zone-$A$. Since $R^{\prime}\gg 2M$, zone-$(1\!\cup\!A)$ is much larger than the black hole interior zone-1. Moreover, most of the zone-$(1\!\cup\!A)$ is outside of the black hole. Thus the zone-$(1\!\cup\!A)$ contains a plenty of space for storing entropy, and most of it is not hidden behind the horizon. Since the “particles” in zone-$(1\!\cup\!A)$ are not physical particles, the physical particle creation in our model can be considered to occur at the sphere of radius $R^{\prime}\gg 2M$  [^1] This picture, of course, should be taken with a grain of salt, because in a more realistic model there would be no sharp boundary between the two zones.. The entanglement entropy of particles created at $R^{\prime}$ can easily be explained through entanglement with degrees of freedom in zone-$(1\!\cup\!A)$, which, in principle, resolves the black hole information paradox.
@@
Then we assume that, in zone-$(1\!\cup\!A)$, there is an over-complete basis of quasi-classical coherent states $|Z_{1\cup A}\rangle$ for both gravitational and matter fields, $|Z_{1\cup A}\rangle=|Z_{\rm gravitational},Z_{\rm matter}\rangle$, obeying
@@
Without a full theory of quantum gravity, we cannot specify the coherent states $|Z_{1\cup A}\rangle$ explicitly. Nevertheless, their essential properties can easily be inferred by heuristic arguments. It is natural to expect that a typical coherent state resembles a classical configuration of gravitational and matter fields. Hence, a typical coherent state behaves approximately classically. In other words, it behaves as a classical hair in zone-$(1\!\cup\!A)$. But we know from classical no hair theorems hawking-ellis ; frolov-novikov that classical hair soon gets radiated away, in terms of gravitational (and matter) waves. Thus our approach predicts that there are two kinds of radiation from the black hole. First there are Hawking particles created in zone-$B$, and second there is a quasi-classical radiation of gravitational and matter waves created in zone-$A$. The two kinds of radiation are entangled with each other, which resolves the black hole information paradox.
@@
Such resolution of the paradox does not depend on any details of quantum gravity and does not involve any exotic phenomena that would contradict general expectations from semiclassical gravity. In particular, a freely falling observer near the horizon will not observe anything out of the ordinary. Indeed, since Hawking particles are not physical near the horizon, such a freely falling observer near the horizon will not observe any Hawking particles.
@@
Appendix A The effect of measurement on the quantum state
In this Appendix we present the basic scheme describing how measurement affects the quantum state. For illustration purposes we describe it for the case of position measurement of electron in the hydrogen atom, but essentially the same scheme works also for measurements in semiclassical and quantum gravity.
To give an operational meaning to the mixed state (3), the position of particle-$B$ must actually be measured by a macroscopic measuring apparatus M. Hence the state (2) generalizes to
$|\Psi\rangle\propto\int d^{3}x\,\psi_{0}({\bf x})|{\bf x}\rangle_{A}|{\bf x}\rangle_{B}|{\rm M}_{\bf x}\rangle,$ (30)
where $|{\rm M}_{\bf x}\rangle$ is the state of the measuring apparatus M corresponding to the measurement outcome ${\bf x}$. So, if the measurement outcome of a single measurement turns out to be ${\bf x}$, we have an effective measurement-induced “collapse”
$|\Psi\rangle\to|{\bf x}\rangle_{A}|{\bf x}\rangle_{B}|{\rm M}_{\bf x}\rangle.$ (31)
Hence the electron is in the state $|{\bf x}\rangle_{A}$, namely, in a state with a well defined position different from the ground state $|\psi_{0}\rangle_{A}$. The state $|{\bf x}\rangle_{A}$ is not stationary, so the electron soon settles down into the ground state and produces electromagnetic radiation during this process. The state of radiation depends on the pre-radiation state $|{\bf x}\rangle_{A}$, so radiation is correlated with particle-$B$ in the state $|{\bf x}\rangle_{B}$.
Appendix B Bogoliubov coefficients
@@
are the Bogoliubov coefficients. For the same reason, the scalar products $(f_{2k},f_{Ak}^{*})$ and $(f_{2k},f_{Bk}^{*})$, i.e., the $\beta$-coefficients in (B), are also negligible, so (35) reduces to